OpenStudy (ammarah):

Graph the function. Label the vertex and axis of symmetry. 1.) y= -2(x+2)(x+6)

OpenStudy (ammarah):

@.Sam.

OpenStudy (ammarah):

the intercept form is y=a(x-p)(x-q)

sam (.sam.):

You should expand it, then you got ax^2+bx+c. Find the vertex using \[Vertex=-\frac {b}{2a}\]

OpenStudy (ammarah):

and x=P+q/2

sam (.sam.):

Are you sure?

ganeshie8 (ganeshie8):

y=a(x-p)(x-q) may be we can simply use the symmetry here x-coordinate of vertex = (p+q)/2 y-coordinate = f((p+q)/2)

OpenStudy (ammarah):

yes so i solved it but my answer truned out to be very huge....can u help me???

ganeshie8 (ganeshie8):

did u change it it :- y= -2(x+2)(x+6) y= -2(x--2)(x--6) p = -2 q = -6

ganeshie8 (ganeshie8):

wat did u get for x-coordinate of vertex ?

OpenStudy (ammarah):

yeah i got 4

ganeshie8 (ganeshie8):

there u went wrong ! u should get -4 negative 4

OpenStudy (ammarah):

how? its -8/-2

ganeshie8 (ganeshie8):

y= -2(x+2)(x+6) y= -2(x--2)(x--6) p = -2 q = -6 x-coordinate of vertex = p+q /2 = -2-6/2 = -8/2 = -4

OpenStudy (ammarah):

wait but its -2

ganeshie8 (ganeshie8):

who said -2 ?

OpenStudy (ammarah):

ohhh im looking at the equationnnnnnnn

ganeshie8 (ganeshie8):

hmm, are we in same page yet ? :)

OpenStudy (ammarah):

yes okay now plug in the x into the equation?

ganeshie8 (ganeshie8):

yup ! plugin x = -4 into the given equaiton, that should give us y-coordinate of vertex

OpenStudy (ammarah):

okay can u help me another one.... y+2x^2 -5x+3

ganeshie8 (ganeshie8):

u mean y = 2x^2 -5x+3

ganeshie8 (ganeshie8):

use the formula given by @.Sam. - first reply

OpenStudy (ammarah):

yeah umm okay

OpenStudy (ammarah):

for x i got -5/4

ganeshie8 (ganeshie8):

careful labout signs !

ganeshie8 (ganeshie8):

y = 2x^2 -5x+3 a = 2 b = -5 c = 3

ganeshie8 (ganeshie8):

x-coordinate of vertex = -b/2a = --5/2*2 = 5/4

OpenStudy (ammarah):

oh yeah

OpenStudy (ammarah):

wait how do u do 5/4^2?

ganeshie8 (ganeshie8):

dint get u

OpenStudy (ammarah):

when u plug it in y how do u solve it?

ganeshie8 (ganeshie8):

y = 2x^2 -5x+3 put x = 5/4 and simplify ?

OpenStudy (ammarah):

i know thats what im having oroblems with

OpenStudy (ammarah):

ohhokay i have a facotring problem 6x^2-9x+3

OpenStudy (ammarah):

so i take out 3?

ganeshie8 (ganeshie8):

thats a good start

OpenStudy (ammarah):

i got 3(2x^2-3x+1)

ganeshie8 (ganeshie8):

2x^2-3x+1 -2*-1 = 2 -2-1 = -3

ganeshie8 (ganeshie8):

those are the magic numbers -2 and -1

ganeshie8 (ganeshie8):

3(2x^2-3x+1) 3(2x^2-2x-1x+1)

ganeshie8 (ganeshie8):

nothing fancy there, i just wrote -3x as -2x-1x

ganeshie8 (ganeshie8):

fine so far ?

OpenStudy (ammarah):

yes i got that too

ganeshie8 (ganeshie8):

good :) 3(2x^2-3x+1) 3(2x^2-2x-1x+1) ^ ^

ganeshie8 (ganeshie8):

factor 2x from them

ganeshie8 (ganeshie8):

3(2x^2-3x+1) 3(2x^2-2x-1x+1) ^ ^ 3(2x(x-1)-1x+1)

OpenStudy (ammarah):

ok then factor -1x+1?

ganeshie8 (ganeshie8):

3(2x^2-3x+1) 3(2x^2-2x-1x+1) ^ ^ 3( 2x(x-1)-1(x-1) )

ganeshie8 (ganeshie8):

yes, u still ok wid above ?

OpenStudy (ammarah):

yes

ganeshie8 (ganeshie8):

3(2x^2-3x+1) 3(2x^2-2x-1x+1) ^ ^ 3( 2x(x-1)-1(x-1) ) ^^^ ^^^

ganeshie8 (ganeshie8):

x-1 is common, so take that out

ganeshie8 (ganeshie8):

3(2x^2-3x+1) 3(2x^2-2x-1x+1) ^ ^ 3( 2x(x-1)-1(x-1) ) ^^^ ^^^ 3( (x-1)(2x-1) )

OpenStudy (ammarah):

I have One More!!!!! Find the value of maxiuximum or minimum

OpenStudy (ammarah):

y=-(x-1)(x+4)

ganeshie8 (ganeshie8):

this thread is lagging for me... very slow mind closing this and asking a new q ? :) pls

OpenStudy (ammarah):

I got for muy answer (-3/2,0 is that right?

ganeshie8 (ganeshie8):

let me check quick

OpenStudy (ammarah):

ok ill post a neq question...

ganeshie8 (ganeshie8):

its leading coefficient is NEGATIVE that means the curve is facing DOWN that means it has a MAXIMUM value

ganeshie8 (ganeshie8):

you need to find the MAXIMUM value

OpenStudy (ammarah):

yes but its ays to fine the point of the vertex

ganeshie8 (ganeshie8):

MAXIMUM value occurs at vertex

OpenStudy (ammarah):

yeah

ganeshie8 (ganeshie8):

y=-(x-1)(x+4) x coordinate = p+q/2 = ?

OpenStudy (ammarah):

-3/2

ganeshie8 (ganeshie8):

yup ! thats x-coordinate where the MAX value occurs. now find the MAX value

ganeshie8 (ganeshie8):

find f(-3/2)

ganeshie8 (ganeshie8):

the y-coordinate of vertex

OpenStudy (ammarah):

and then when i plugged it in i got 0 is that rigfht i hate solving fractions.....

ganeshie8 (ganeshie8):

i hate too...

ganeshie8 (ganeshie8):

http://www.wolframalpha.com/input/?i=y%3D-%28x-1%29%28x%2B4%29%2C+x+%3D+-3%2F2 check where u did wrong

OpenStudy (ammarah):

like sometimes i et stuck when i have to subract them.,

ganeshie8 (ganeshie8):

its okay, takes practice id guess

ganeshie8 (ganeshie8):

adding fractions is not the point here, finding the max value ok... just see that MAX value occurs at VERTEX

OpenStudy (ammarah):

i know but when -3/2 is pkugged in i dk whow to solve after that......

ganeshie8 (ganeshie8):